OpenStudy (annipuppi):

Can you help me with these 3 compound inequalities? They're probably really easy, but I don't remember how to solve them.

OpenStudy (annipuppi):

-2t+2<4 and 2t<6 3x-1<=5 or 2x-4>=x -3t<=12 and -2t>-6

OpenStudy (annipuppi):

@robtobey I really only need help with the first one if that's all you want to help with